OpenStudy (anonymous):

Find the period of the function. y = 5 cos 1/2 x

OpenStudy (anonymous):

period of \[\cos(bx)\] is \[\frac{2\pi}{b}\] in your example \(b=\frac{1}{2}\)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

so period is \[\frac{2\pi}{\frac{1}{2}}=2\pi\times 2=4\pi\]

OpenStudy (anonymous):

ok thank you! can u help me with something else?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 17 feet and the height from the ground to the front doors is 7 feet, how long is the ramp? (Round to the nearest tenth.)

OpenStudy (anonymous):



OpenStudy (anonymous):

find the hypotenuse via pythagoras \[h=\sqrt{7^2+17^2}\]

OpenStudy (anonymous):

18.4?